# Organizing Unstructured Image Collections using Natural Language

### v4

@@ Appendix I Failure Case Analysis
We evaluate this on two image collections containing fine-grained criteria: CUB200 (Wah et al., 2011) and Stanford Cars196 (Khosla et al., 2011). Our framework successfully discovers the fine-grained criteria Bird species for CUB200 and Car model for Cars196. As shown in Tab. 34, when uncovering fine-grained substructures, integrating the FineR prompting strategy significantly improves performance by up to +15.0% CAcc and +12.2% SAcc, achieving results comparable to FineR itself. This demonstrates the flexibility of our system, allowing future adaptations to specific application needs, such as fine-grained image collections.
Appendix P Further Details of the Application Study
In this section, we present additional implementation details, evaluation results, and findings for the application study discussed in § 6 of the main paper. Specifically, § P.3 offers further evaluation results and implementation details on using our predicted distribution to train a debiased model with GroupDRO (Sagawa et al., 2020). § P.1 outlines the implementation of the user study that assesses the alignment between predicted biases and human judgments, along with comprehensive findings for all studied occupations and identified criteria. Finally, § P.2 provides additional insights from the analysis of social media image popularity.
P.1 Further Details on Discovering Novel Bias in Text-to-Image Diffusion Models
@@
Image Generation for the Subject Occupation: Following prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016), we selected nine occupations for our study: three stereotypically biased towards females (Nurse, Cleaning staff, Call center employee), three biased towards males (CEO, Firefighter, Basketball player), and three considered gender-neutral (Teacher, Computer user, Marketing coordinator). We used two state-of-the-art T2I diffusion model, DALL·E3 (Betker et al., 2023) and Stable Diffusion (SDXL) (Podell et al., 2024) to generate 100 images for each occupation for our study. This resulted in a total of 1,800 images. For each occupation, we provide some examples of images generated by DALL·E3 in Figure 22, while provide some examples of images generated by SDXL in Figure 23. We only used the simple prompt “A portrait photo of a $<$OCCUPATION$>$” for image generation for all occupations and did not include any potential biases in the prompt.
Bias Discovery and Quantification: We applied our method to 1,800 generated images and automatically identified 10 grouping criteria (bias dimensions) along with their predicted distributions for each occupation image set. For this study, we utilized the mid-granularity output of our system. To evaluate the biases, we first identified the dominant cluster for each criterion—the cluster containing the largest number of images—as the bias direction. We then calculated the normalized entropy of the distribution for each criterion of the occupation’s images to determine the bias intensity score, following the method proposed by D’Incà et al. (2024):
@@
The Absolute Mean Error (AME) between the bias intensity scores predicted by our system and those rated by humans (scaled to 0 to 1) was 0.1396. Additionally, our system’s predicted bias directions aligned with human evaluations 72.3% of the time, with most discrepancies occurring in the criteria of “Age group,” “Skin tone,” and “Accessories worn.” These findings indicate a strong correlation between our system’s predictions and human judgments, validating the effectiveness of our approach. Detailed user study results are provided in § P.1. We believe the discrepancies in certain criteria may be due to the influence of personal subjective cognition on respondents’ answers. In Figure 19, we present the human evaluation results, averaged across all participants for each model, occupation, and criterion, with the human ratings scaled from 0 to 1.
@@
We speculate that DALL·E3’s superior performance in mitigating well-known biases may be attributed to its “guardrails” (OpenAI, 2022b), designed as part of its industrial deployment to avoid amplifying social biases via its easily accessible APIs. However, these guardrails do not prevent it from exhibiting biases along the novel dimensions discovered by our method, as these dimensions remain understudied. This observation highlights the importance of studying novel biases that could potentially exist in widely used T2I generative models to prevent further bias amplification.
P.2 Further Details on Analyzing Social Media Image Popularity
@@
With the rise of image-centric content on social media platforms like Instagram, Flickr, and TikTok, understanding what makes an image popular has become crucial for applications such as marketing, content curation, and recommendation systems. Traditional research often approaches image popularity as a regression problem (Ortis et al., 2019; Cheng et al., 2024), utilizing metadata like hashtags, titles, or follower counts. However, the specific semantic visual elements that contribute to an image’s popularity remain largely unexplored. In this study, we applied our proposed method to automatically categorize social media images based on semantic visual elements across different dimensions (criteria). By analyzing these interpretable results alongside image popularity metrics (e.g., number of views), we gained insights into the factors contributing to virality and identified common visual traits among popular images. These insights can provide valuable guidance for content creators and advertisers, enhancing productivity and informing strategic decision-making.
@@
To expand on the discussion in § 6 of the main paper, we present the complete findings across all ten discovered criteria in Figure 20. Notably, we consistently observed a sharp semantic contrast between the visual elements in top trending images and those in the mainstream images across all ten criteria. For instance, there is a contrast between Urban sophisticated and Modern minimalist under Interior Design, Rustic architecture and Modern architecture under Architecture Style, and Event venues versus Urban residential areas under Location.
@@
This recurring observation reinforces the idea that viral (or trending) content tends to capture more attention, likely because it features novel, surprising, or striking visual elements. Humans are inherently attracted to stimuli that deviate from the norm (Priester et al., 2004; Bruni et al., 2012; Palmer and Gore, 2014). On the other hand, widely uploaded yet “neutral” content is shared more often due to its familiarity and broad appeal, though it is less likely to provoke the strong emotional responses that fuel virality. We believe the insights generated by our method could offer valuable guidance to social media platform practitioners, helping them tailor their content more effectively to target audiences and gain a deeper understanding of social media image trends from various perspectives.
P.3 Confirming and Mitigating Dataset Bias
@@
Confirming and Mitigating Dataset Bias: Given an image collection that contains spurious correlations (Geirhos et al., 2020), we are curious whether we can proactively find this issue caused by data bias directly from the training images without relying on either the annotations (Sagawa et al., 2020) or post hoc misclassified images (Kim et al., 2024). As a case study, we applied the proposed $\mathcal{X}$-Cluster framework to the 162k training images of the CelebA (Liu et al., 2015) dataset—a binary hair color classification dataset where the target label “Blond” is spuriously correlated with the demographic attribute “Female” in its training split.
@@
Findings: As expected, our method identified the grouping criteria Hair color and Gender. Next, we analyzed the predicted gender distributions within the “Blond” and “Not Blond” (all other colors) clusters. As shown in Figure 21, we observed that the gender distribution within the “Blond” cluster is highly skewed, with 86.5% of the images representing females, closely matching the ground-truth distribution (94.3%). Such an imbalance confirms the potential issue of spurious correlations between “Blond” and “Female”. To further validate this observation, following B2T (Kim et al., 2024), we used our predicted distributions to train a debiased model with GroupDRO (Sagawa et al., 2020) and compared it with other unsupervised bias mitigation methods, including JTT (Liu et al., 2021), CNC (Zhang et al., 2022), B2T, and GroupDRO trained with ground-truth labels. As shown in Tab. 35, our debiased model achieved robust performance, comparable to that of B2T, demonstrating the reliability of its discovered distributions.
Additional Evaluation: To further evaluate the prediction quality of our method for hair color and gender, we used the ground-truth labels from the CelebA dataset (Liu et al., 2015) to assess the classification accuracy of them. Our method achieved an impressive classification accuracy of 99.1% for gender and 87.4% for hair color on the 162,770 training images, demonstrating its effectiveness for uncovering gender and hair color substructures within the training set.
In addition, we quantified the spurious correlation between hair color and gender using the metric proposed by Yang et al. (2023). Specifically, given the correlated gender attribute distribution $A$ and the target hair color distribution $Y$, we computed the normalized mutual information between $A$ and $Y$ to quantify the spurious correlation as:
@@
Implementation Details of Training GroupDRO: To conduct debiased training using GroupDRO (Sagawa et al., 2020), we first used our predicted distribution to define four distinct training groups, rather than relying on the ground-truth distribution. We closely followed the training protocol outlined in B2T (Kim et al., 2024) and GroupDRO (Sagawa et al., 2020). Specifically, we fine-tuned a ResNet-50 (He et al., 2016) model pre-trained on ImageNet (Deng et al., 2009), using the training split of the CelebA dataset (Liu et al., 2015). The training was performed using the SGD optimizer (Ruder, 2016) with a momentum of 0.9, a batch size of 64, and a learning rate of $1\times 10^{-5}$. We applied a weight decay of 0.1 and set the group adjustment parameter to zero. The model was trained over 50 epochs. For evaluation, we reported both the average and worst-group test accuracies, selecting the model from the epoch that achieves the highest worst-group accuracy on the validation set. The final evaluation and comparison results are provided in Tab. 35.
Appendix Q Why LLMs Improve Image Clustering?
@@
The most compelling aspect of this work lies in our $\mathcal{X}$-Cluster framework’s ability to transform large volumes of unstructured images into natural language and leverage the advanced text understanding and summarization capabilities of LLMs to tackle the challenging Open-ended Semantic Multiple Clustering (OpenSMC) task. This approach draws inspiration from the use of LLMs in the Topic Discovery task within the NLP domain (Eklund and Forsman, 2022). Our core motivation is: “If LLMs can discover topics from documents and organize them, then by converting images into text, we can similarly use LLMs to organize unstructured images.”
@@
Traditional clustering methods (Estivill-Castro, 2002; Caron et al., 2018; Van Gansbeke et al., 2020; Li et al., 2024; Yu et al., 2024) often depend on pre-defined criteria, pre-determined numbers of clusters, fixed feature representations (which require training), and are typically not interpretable. These limitations hinder their applicability to diverse datasets in open-world scenarios, as they demand significant human priors and retraining for each new dataset.
@@
In contrast, LLMs (OpenAI, 2022a, 2023; Touvron et al., 2023; Meta, 2024b, a) excel at understanding, summarizing, and reasoning over high-level semantics expressed in natural language across diverse domains (e.g., everyday content, cultural knowledge, or medical content). Operating in a zero-shot (Kojima et al., 2022), interpretable manner, LLMs are uniquely suited to the SMC task, which aims to discover meaningful and interpretable clustering criteria without requiring prior knowledge or training. By integrating LLMs with MLLMs (Liu et al., 2024b) into the carefully designed $\mathcal{X}$-Cluster framework, we enable the discovery and refinement of clustering criteria directly from the dataset’s content, followed by automatic grouping of the dataset. This design allows our framework to overcome the rigid assumptions of traditional clustering methods, making it automatic, generalizable, and training-free. Our approach provides a novel perspective, demonstrating how clustering tasks can evolve beyond traditional paradigms.
@@
Challenges of employing LLMs to facilitate the SMC task. The main challenge of employing LLMs for the SMC task lies in accurately translating visual content from images into natural language that LLMs can effectively reason with. This is evident from the sensitivity analysis results in App. N: $\mathcal{X}$-Cluster’s performance improves with larger or more powerful MLLMs (see Figure 18 (a)), while it remains relatively insensitive to the specific choice of LLM (see Figure 18 (b)). In other words, the quality of image captions generated by MLLMs is critical for the effective use of LLMs in the SMC task. Specifically, in the first stage of $\mathcal{X}$-Cluster (criteria proposal), captions need to be as comprehensive as possible to provide rich information for LLMs to discover grouping criteria. In the second stage (semantic grouping), criterion-specific captions should precisely capture relevant visual content to provide accurate information for assigning images to clusters.
@@
To enhance caption quality, techniques such as MLLM model ensembling, prompt ensembling (Liu et al., 2024c), or stronger models like GPT-4V (OpenAI, 2023) can improve comprehensiveness. For better precision, advanced prompting methods like CoT (Wei et al., 2022) or FineR (Liu et al., 2024e) can capture nuanced details, while hallucination mitigation tools like Visual Fact Checker (Ge et al., 2024) can reduce noise caused by hallucinations. However, these techniques increase computational costs and framework complexity. In this work, we choose to keep $\mathcal{X}$-Cluster simple yet effective, and we outline these potential improvements for future practitioners.
Appendix R Future Work
@@
Closed-Loop Optimization. In this work, we designed our prompts following the Iterative Prompt Engineering methodology (DeepLearning.AI, 2024) introduced by Isa Fulford and Andrew Ng. In App. E, we provide the exact LLM and MLLM prompts used in our framework and break down each prompt to explain the objectives and purposes behind each design choice. These explanations cover elements such as system prompts, input formatting, task and sub-task instructions, and output instructions. Our focus in this work is on creating a highly generalizable framework, $\mathcal{X}$-Cluster, and we do not perform any closed-loop, dataset-specific prompt optimizations. However, in future work or application scenarios where a labeled training/validation dataset is available, practitioners could build upon our design objectives. By leveraging our proposed evaluation metrics (see § 5) for each step, it would be possible to develop a Open-ended Semantic Multiple Clustering (OpenSMC) system with a closed-loop optimization pipeline to achieve improved performance.
@@
  • Audio Data: Speech-to-Text models like Whisper (Radford et al., 2023) can convert audio data into text, enabling subsequent analysis with $\mathcal{X}$-Cluster.
  • Tabular Data: Table-to-Text models, such as TabT5 (Andrejczuk et al., 2022), can translate tabular data into text, making it compatible with $\mathcal{X}$-Cluster. For tables containing figures, modern MLLMs like LLaVA-Next, which support both OCR and image-to-text capabilities, can handle these elements to create a unified textual representation for $\mathcal{X}$-Cluster.
  • Protein Structures: Protein structure-to-text models, such as ProtChatGPT (Wang et al., 2024a), can convert protein sequences into textual descriptions for analysis with $\mathcal{X}$-Cluster.
  • Point Cloud Data: 3D captioning models, like Cap3D (Luo et al., 2024a), can transform point cloud data or rendered 3D models into text, enabling their analysis using $\mathcal{X}$-Cluster.
@@
Figure 22: Samples of DALL·E3 generated images. For each occupation, the simple prompt “A portrait photo of a $<$OCCUPATION$>$”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to DALL·E3 to generate 100 images. We present a random sample of 30 generated images.
Figure 23: Samples of SDXL generated images. For each occupation, the simple prompt “A portrait photo of a $<$OCCUPATION$>$”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to SDXL to generate 100 images. We present a random sample of 30 generated images.

### v5

@@ Appendix I Qualitative Analysis
We evaluate this on two image collections containing fine-grained criteria: CUB200 [wah2011caltech] and Stanford Cars196 [khosla2011novel]. Our framework successfully discovers the fine-grained criteria Bird species for CUB200 and Car model for Cars196. As shown in Tab. 32, when uncovering fine-grained substructures, integrating the FineR prompting strategy significantly improves performance by up to +15.0% CAcc and +12.2% SAcc, achieving results comparable to FineR itself. This demonstrates the flexibility of our system, allowing future adaptations to specific application needs, such as fine-grained image collections.
Appendix Q Further Details of the Application Study
In this section, we present additional implementation details, evaluation results, and findings for the application study discussed in Sec. 6 of the main paper. Specifically, § Q.3 offers further evaluation results and implementation details on using our predicted distribution to train a debiased model with GroupDRO [sagawa2019distributionally]. § Q.1 outlines the implementation of the user study that assesses the alignment between predicted biases and human judgments, along with comprehensive findings for all studied occupations and identified criteria. Finally, § Q.2 provides additional insights from the analysis of social media image popularity.
Q.1 Further Details on Discovering Novel Bias in Text-to-Image Diffusion Models
@@
Image Generation for the Subject Occupation: Following prior studies [bianchi2023easily, bolukbasi2016man], we selected nine occupations for our study: three stereotypically biased towards females (Nurse, Cleaning staff, Call center employee), three biased towards males (CEO, Firefighter, Basketball player), and three considered gender-neutral (Teacher, Computer user, Marketing coordinator). We used two state-of-the-art T2I diffusion model, DALL·E3 [betker2023improving] and Stable Diffusion (SDXL) [podell2023sdxl] to generate 100 images for each occupation for our study. This resulted in a total of 1,800 images. For each occupation, we provide some examples of images generated by DALL·E3 in Figure 14, while provide some examples of images generated by SDXL in Figure 15. We only used the simple prompt “A portrait photo of a <<OCCUPATION>>” for image generation for all occupations and did not include any potential biases in the prompt.
Bias Discovery and Quantification: We applied our method to 1,800 generated images and automatically identified 10 grouping criteria (bias dimensions) along with their predicted distributions for each occupation image set. For this study, we utilized the mid-granularity output of our system. To evaluate the biases, we first identified the dominant cluster for each criterion—the cluster containing the largest number of images—as the bias direction. We then calculated the normalized entropy of the distribution for each criterion of the occupation’s images to determine the bias intensity score, following the method proposed by d2024openbias:
@@
The Absolute Mean Error (AME) between the bias intensity scores predicted by our system and those rated by humans (scaled to 0 to 1) was 0.1396. Additionally, our system’s predicted bias directions aligned with human evaluations 72.3% of the time, with most discrepancies occurring in the criteria of “Age group,” “Skin tone,” and “Accessories worn.” These findings indicate a strong correlation between our system’s predictions and human judgments, validating the effectiveness of our approach. Detailed user study results are provided in § Q.1. We believe the discrepancies in certain criteria may be due to the influence of personal subjective cognition on respondents’ answers. In Figure 11, we present the human evaluation results, averaged across all participants for each model, occupation, and criterion, with the human ratings scaled from 0 to 1.
@@
We speculate that DALL·E3’s superior performance in mitigating well-known biases may be attributed to its “guardrails” [dalle_guards], designed as part of its industrial deployment to avoid amplifying social biases via its easily accessible APIs. However, these guardrails do not prevent it from exhibiting biases along the novel dimensions discovered by our method, as these dimensions remain understudied. This observation highlights the importance of studying novel biases that could potentially exist in widely used T2I generative models to prevent further bias amplification.
Q.2 Further Details on Analyzing Social Media Image Popularity
@@
With the rise of image-centric content on social media platforms like Instagram, Flickr, and TikTok, understanding what makes an image popular has become crucial for applications such as marketing, content curation, and recommendation systems. Traditional research often approaches image popularity as a regression problem [ortis2019prediction, cheng2024retrieval], utilizing metadata like hashtags, titles, or follower counts. However, the specific semantic visual elements that contribute to an image’s popularity remain largely unexplored. In this study, we applied our proposed method to automatically categorize social media images based on semantic visual elements across different dimensions (criteria). By analyzing these interpretable results alongside image popularity metrics (e.g., number of views), we gained insights into the factors contributing to virality and identified common visual traits among popular images. These insights can provide valuable guidance for content creators and advertisers, enhancing productivity and informing strategic decision-making.
@@
To expand on the discussion in Sec. 6 of the main paper, we present the complete findings across all ten discovered criteria in Figure 12. Notably, we consistently observed a sharp semantic contrast between the visual elements in top trending images and those in the mainstream images across all ten criteria. For instance, there is a contrast between Urban sophisticated and Modern minimalist under Interior Design, Rustic architecture and Modern architecture under Architecture Style, and Event venues versus Urban residential areas under Location.
@@
This recurring observation reinforces the idea that viral (or trending) content tends to capture more attention, likely because it features novel, surprising, or striking visual elements. Humans are inherently attracted to stimuli that deviate from the norm [priester2004and, bruni2012role, palmer2014theory]. On the other hand, widely uploaded yet “neutral” content is shared more often due to its familiarity and broad appeal, though it is less likely to provoke the strong emotional responses that fuel virality. We believe the insights generated by our method could offer valuable guidance to social media platform practitioners, helping them tailor their content more effectively to target audiences and gain a deeper understanding of social media image trends from various perspectives.
Q.3 Confirming and Mitigating Dataset Bias
@@
Confirming and Mitigating Dataset Bias: Given an image collection that contains spurious correlations [geirhos2020shortcut], we are curious whether we can proactively find this issue caused by data bias directly from the training images without relying on either the annotations [sagawa2019distributionally] or post hoc misclassified images [kim2024discovering]. As a case study, we applied the proposed $\mathcal{X}$-Cluster framework to the 162k training images of the CelebA [liu2015deep] dataset—a binary hair color classification dataset where the target label “Blond” is spuriously correlated with the demographic attribute “Female” in its training split.
@@
Findings: As expected, our method identified the grouping criteria Hair color and Gender. Next, we analyzed the predicted gender distributions within the “Blond” and “Not Blond” (all other colors) clusters. As shown in Figure 13, we observed that the gender distribution within the “Blond” cluster is highly skewed, with 86.5% of the images representing females, closely matching the ground-truth distribution (94.3%). Such an imbalance confirms the potential issue of spurious correlations between “Blond” and “Female”. To further validate this observation, following B2T [kim2024discovering], we used our predicted distributions to train a debiased model with GroupDRO [sagawa2019distributionally] and compared it with other unsupervised bias mitigation methods, including JTT [liu2021just], CNC [zhang2022correct], B2T, and GroupDRO trained with ground-truth labels. As shown in Tab. 33, our debiased model achieved robust performance, comparable to that of B2T, demonstrating the reliability of its discovered distributions.
Additional Evaluation: To further evaluate the prediction quality of our method for hair color and gender, we used the ground-truth labels from the CelebA dataset [liu2015deep] to assess the classification accuracy of them. Our method achieved an impressive classification accuracy of 99.1% for gender and 87.4% for hair color on the 162,770 training images, demonstrating its effectiveness for uncovering gender and hair color substructures within the training set.
In addition, we quantified the spurious correlation between hair color and gender using the metric proposed by yang2023change. Specifically, given the correlated gender attribute distribution $A$ and the target hair color distribution $Y$, we computed the normalized mutual information between $A$ and $Y$ to quantify the spurious correlation as:
@@
Implementation Details of Training GroupDRO: To conduct debiased training using GroupDRO [sagawa2019distributionally], we first used our predicted distribution to define four distinct training groups, rather than relying on the ground-truth distribution. We closely followed the training protocol outlined in B2T [kim2024discovering] and GroupDRO [sagawa2019distributionally]. Specifically, we fine-tuned a ResNet-50 [he2016deep] model pre-trained on ImageNet [deng2009imagenet], using the training split of the CelebA dataset [liu2015deep]. The training was performed using the SGD optimizer [ruder2016overview] with a momentum of 0.9, a batch size of 64, and a learning rate of $1\times 10^{-5}$. We applied a weight decay of 0.1 and set the group adjustment parameter to zero. The model was trained over 50 epochs. For evaluation, we reported both the average and worst-group test accuracies, selecting the model from the epoch that achieves the highest worst-group accuracy on the validation set. The final evaluation and comparison results are provided in Tab. 33.
Appendix R Why LLMs Improve Image Clustering?
@@
The most compelling aspect of this work lies in our $\mathcal{X}$-Cluster framework’s ability to transform large volumes of unstructured images into natural language and leverage the advanced text understanding and summarization capabilities of LLMs to tackle the challenging Open-ended Semantic Multiple Clustering (OpenSMC) task. This approach draws inspiration from the use of LLMs in the Topic Discovery task within the NLP domain [eklund2022topic]. Our core motivation is: “If LLMs can discover topics from documents and organize them, then by converting images into text, we can similarly use LLMs to organize unstructured images.”
@@
Traditional clustering methods [estivill2002so, caron2018deep, van2020scan, li2023image, yu2024multiple] often depend on pre-defined criteria, pre-determined numbers of clusters, fixed feature representations (which require training), and are typically not interpretable. These limitations hinder their applicability to diverse datasets in open-world scenarios, as they demand significant human priors and retraining for each new dataset.
@@
In contrast, LLMs [chatgpt, openai2023gpt4, touvron2023llama, meta2024llama3, meta2024llama31] excel at understanding, summarizing, and reasoning over high-level semantics expressed in natural language across diverse domains (e.g., everyday content, cultural knowledge, or medical content). Operating in a zero-shot [kojima2022large], interpretable manner, LLMs are uniquely suited to the SMC task, which aims to discover meaningful and interpretable clustering criteria without requiring prior knowledge or training. By integrating LLMs with MLLMs [liu2024llava] into the carefully designed $\mathcal{X}$-Cluster framework, we enable the discovery and refinement of clustering criteria directly from the dataset’s content, followed by automatic grouping of the dataset. This design allows our framework to overcome the rigid assumptions of traditional clustering methods, making it automatic, generalizable, and training-free. Our approach provides a novel perspective, demonstrating how clustering tasks can evolve beyond traditional paradigms.
@@
Challenges of employing LLMs to facilitate the SMC task. The main challenge of employing LLMs for the SMC task lies in accurately translating visual content from images into natural language that LLMs can effectively reason with. This is evident from the sensitivity analysis results in Supp. O: $\mathcal{X}$-Cluster’s performance improves with larger or more powerful MLLMs (see Figure 10 (a)), while it remains relatively insensitive to the specific choice of LLM (see Figure 10 (b)). In other words, the quality of image captions generated by MLLMs is critical for the effective use of LLMs in the SMC task. Specifically, in the first stage of $\mathcal{X}$-Cluster (criteria proposal), captions need to be as comprehensive as possible to provide rich information for LLMs to discover grouping criteria. In the second stage (semantic grouping), criterion-specific captions should precisely capture relevant visual content to provide accurate information for assigning images to clusters.
@@
To enhance caption quality, techniques such as MLLM model ensembling, prompt ensembling [liu2024visual], or stronger models like GPT-4V [openai2023gpt4] can improve comprehensiveness. For better precision, advanced prompting methods like CoT [wei2022chain] or FineR [liu2024democratizing] can capture nuanced details, while hallucination mitigation tools like Visual Fact Checker [ge2024visual] can reduce noise caused by hallucinations. However, these techniques increase computational costs and framework complexity. In this work, we choose to keep $\mathcal{X}$-Cluster simple yet effective, and we outline these potential improvements for future practitioners.
Appendix S Future Work
@@
Closed-Loop Optimization. In this work, we designed our prompts following the Iterative Prompt Engineering methodology [deeplearningai_2024] introduced by Isa Fulford and Andrew Ng. In Supp. F, we provide the exact LLM and MLLM prompts used in our framework and break down each prompt to explain the objectives and purposes behind each design choice. These explanations cover elements such as system prompts, input formatting, task and sub-task instructions, and output instructions. Our focus in this work is on creating a highly generalizable framework, $\mathcal{X}$-Cluster, and we do not perform any closed-loop, dataset-specific prompt optimizations. However, in future work or application scenarios where a labeled training/validation dataset is available, practitioners could build upon our design objectives. By leveraging our proposed evaluation metrics (see LABEL:sec:method_assessment) for each step, it would be possible to develop a Open-ended Semantic Multiple Clustering (OpenSMC) system with a closed-loop optimization pipeline to achieve improved performance.
@@
  • Audio Data: Speech-to-Text models like Whisper [radford2023robust] can convert audio data into text, enabling subsequent analysis with $\mathcal{X}$-Cluster.
  • Tabular Data: Table-to-Text models, such as TabT5 [andrejczuk2022table], can translate tabular data into text, making it compatible with $\mathcal{X}$-Cluster. For tables containing figures, modern MLLMs like LLaVA-Next, which support both OCR and image-to-text capabilities, can handle these elements to create a unified textual representation for $\mathcal{X}$-Cluster.
  • Protein Structures: Protein structure-to-text models, such as ProtChatGPT [wang2024protchatgpt], can convert protein sequences into textual descriptions for analysis with $\mathcal{X}$-Cluster.
  • Point Cloud Data: 3D captioning models, like Cap3D [luo2024scalable], can transform point cloud data or rendered 3D models into text, enabling their analysis using $\mathcal{X}$-Cluster.
@@
Figure 14: Samples of DALL·E3 generated images. For each occupation, the simple prompt “A portrait photo of a <<OCCUPATION>>”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to DALL·E3 to generate 100 images. We present a random sample of 30 generated images.
Figure 15: Samples of SDXL generated images. For each occupation, the simple prompt “A portrait photo of a <<OCCUPATION>>”, that does not reference any potential bias dimensions such as gender, race or hair color, is fed to SDXL to generate 100 images. We present a random sample of 30 generated images.